Gas systems project

Europe/Zurich
Beatrice Mandelli (CERN)

GAS M&O meeting 29 January 2021

Present: Beatrice, Roberto, Fred, Patrick, LPh, Kacper, Michal, Andrea, Gianluca, Mara

 

ALICE #3104

TPC

To remember:

  • To remove filter in purifier column

From previous meeting (05/10)

  • Krohne: only present one readying for the full system (distribution is just one channel) —> Leonardo gave to Michal the formula (15/10) —> Michal to implement
  • To do fine tuning of pump

TRD

  • difficult to restart: need to work on PID of pump and distribution —> Gianluca
  • check of Xe recuperation on-going
    • different logic wrt logo: in the stand-by pump in logo always ON, with new software not
    • to check if by leaving pump ON all time, recuperation would be better
    • other difference is in the buffer pressure: it wa increase by Chilo during Covid period at ~1.7 bar —> now we’ll decrease to Run 2 value

To remember:

  • T sensor not present in the pump: LPh to check if we can add
  • H2O now it is better (even if still a bit higher, purifier material not changed) to keep checking
  • CO2 didn’t open in mixer: why? —> to check, maybe a leak in Tescom?
  • problem of noise in the read-out of Krohne to investigate
  • Brown bottle to fill: tested in another bottle and from the test we know now that it has to be done all manual --> LPh to check with Chilo when to do
  • disk de rupture: they work —> LPh has to contact the safety

HMP

From previous meeting (03/08):

  • Distribution rack in alarm —> LPh to check

MTR

  • pressure sensor on the chamber not present: LPh to order the sensor (ATEX). Nobody from MTR team at CERN to check where we can put the sensor —> to remember for March
  • no gas need before March

From previous meeting (22/01)

  • to connect dummy chamber electrically —> LPh
  • new software with dummy chamber to discuss —> Michal/Gianluca

To remember

  • leak test and maintenance in the distribution —> LPh/Leonardo
  • No dummy chamber present in software —> Michal

TOF

  • GC analysis: SF6 at 6.7% instead of 7% —> Mara to calculate maximum error of MFC
    • We don’t touch the mixer MFCs
    • MFCs are working at 10% of full range —> Mara to check Run2 date and propose new MFCs

MCH

  • system restarted but flow is 295 l/h instead of 490 l/h —> Lph/Gianluca to investigate

To remember:

  • still some work in PID tuning
  • Maintenance of distribution to do: to check with TOF people if we can do

PGS

  • N2 pressure probably wrong after intervention of EN. Probably they changed the pressure sensor like in ATLAS —> Lph to send an email to David and E Montava with in cc Michal to check about this pressure (and the others)

OTHER

To remember:

  • Analysis chain 3 on, to add calibration line

 

ATLAS #3104

MDT

From previous meeting (22/01)

  • rack 73: to change motherboard and/or ELMB. One at the time
  • some alarms received for some pressure sensors by piquet —> reset to do
    • to also check grounding —> to check with Patrick

To remember:

  • problem in one VAISALA reading but it is not the sensor —> to investigate
  • regulation valve material to buy asap —> Kacper

From previous meeting (20/05):

  • Reset of pressure sensors giving problem (elog #2489)
    • To check the grounding

RPC

  • Leak test on-going for Purifier 2 and 3 with CO2: some leaks found. Repairing on-going
  • Leak test on-going for buffer of the exhaust

From previous meeting:

  • one cable UX-US damaged: Patrick to give to Kacper the reference. Also to check status of other cables

To remember:

  • maintenance purifier: leaks found
    • Purifier 1: leak on the bottom: it needs to be empty —> next year
  • Leonardo to update maintenance procedure

From previous meeting (13/10)

  • Still some electrical problems:
    • pump regulation valve not working correctly: to replace electrical part —> Beatrice to propose a date
    • to test all cables UX-US

TGC

  • some help for the mapping in gas distribution for the new channels

TFC

From previous meeting (22/01)

  • stop caused by a spike in pressure sensor in C-side distribution: not yet clear why it happened

From previous meeting (23/10)

  • Krohne new formula prepared by Leonardo —> Michal to implement
  • Krone exhaust not working properly: to check if it is the krohne or our electrical part. Patrick to try to check the electrical part
  • Kacper to check for regeneration of purifier material

MM

From previous meeting:

  • detection head installed but IR not ordered yet: Beatrice to check with Paolo if they want to order

ID

  • Still waiting a reply for the installation of the IDEP chassis —> Kacper to send an email asap. Now possible from w7

PGS

  • Ar, N2, CO2: alarm blocked for intervention
    • Michal explained that now for ATLAS and LHCb we can put OFF the PCO to not receive the alarm. This has to be done by gas team and not by EN team

 

CMS #3100

RPC

  • increase flow to 500 l/h and put MFC iC4H10 at 4.5% for the February Run
  • 2 IRs recalibrated by LPh but he doesn’t know why there was a drift after one month —> to keep an eye
  • To add in the checklist value of IR, MFC setting and GC measurements
  • Adding volume chambers in WinCCOA on-going by Mara
  • flowcell reading problem: all at 0 l/h: intervention b Michal, who detected a problem for racks 64 and 68 (it already happened last year) —> to investigate if it is the power supply or the canbus cables —> Michal/Andrea
  • Stop of system due to leak detection in 2 racks: after investigation nothing found. Mara to send an email to Nicolas to check if he can tell us what gas triggered the detection head
  • There is a lost of 200 mbar in the pump-purifier line —> Andrea to check
  • Alarm on pump regulation due to increase of flow: the real problem is on the group regulation valves that are almost at 90% —> Andrea/Mara/Roberto to take action

To remember:

  • difficult to reach 1000 l/h. Only possible with one MFC: Kacper/Fabio to check the other MFC
  • 200 l/h is very low wrt the MFC range, so probably we have a big error: to think to buy a set of MFC for lower flows —> Beatrice/Mara

CSC

  • stopped due to detection leak in RPC —> Andrea to contact Nicolas Broca and be sure that problem will be fixed next week

To remember to do

  • need to do purifier maintenance (and change festo block which is leaking): to discuss but for sure after February Run
  • Found leak in safety valve of purifier (~800 ppm): to replace next year
  • To put scale sticker on all bubblers —> Andrea

DT

  • meeting with DT people —> Andrea/Mara to prepare summary for flow problem

To remember

  • Still problem of buffer pressure decreasing when the system stops —> to investigate
  • MFC calibration postponed to w8 or w10
  • Profibus cable has to be change: need to stop the system —> next year

GEM

  • restarted in Ar/CO2 70/30
  • input pressure at 38 mbar for ~150 l/h (for future GEM would like ~ 200 l/h)—> to think to change the pressure sensors with the 0-100 mbar —> to buy the special series
  • safety valves: 50 mbar in + side and 70 mbar in - side. To change in positive side

PGS

  • iC4H10 alarm triggered during power-cut of 28/01 at 12:20 —> to inform David

 

LHCb #3103

RICH1

From previous meeting (08/01):

  • need to change membrane of pump 1

RICH2

  • PLC crash not understood: change of PLC and send it to Schiender —> Michal/Marc
  • still some work to do, list presented by Patrick

To remember

  • strange PID for RICH2 (Kc=25, Ti=20): they were changed on 03/06 at 2pm —> Why?
    • can we go back to old values —> Mara to check with Michal for the logic
  • PT5005 changed but still spikes visible: to think about
  • problem of high injection of CO2 from the mixer
    • probably due to purifier (leak in column A + not saturated in CO2)
    • leak already present in 2018 (elog #2886)

ANALYSIS CHAIN 3

From previous meeting:

  • probably a leak in the electro-valve of Ar-100 ppm H2O —> to change electrovalve
  • new Ar-100 ppm H2O to connect

PGS

  • Leak in the main CF4 line —> Patrick to inform David

GENERAL

  • On Wednesday afternoon stop of systems for the gas detection check (Nicolas Broca)

 

NON-LHC

  • 904 works on-going for RPC pipes —> LPh to show to Andrea, who can take care of restarting the system

 

SOFTWARE

  • New feature: alarms will be divided per detector, so detector people can subscribe.
  • This new feature and alarms/warnings are being implemented in ALICE, in next week the other experiments

From previous meeting

  • New feature for alarm when system is OFF: to provide to Michal the list of alarms we want to have when system are OFF —> Mara/Gianluca with experiment technical responsible by week 5
  • Profibus/canbus monitoring: Michal shows where we have and where it could be useful to have (based on history of problems in the past). To discuss if regeneration of code is worst for these system
  • We need to define how to proceed for the classification of warnings and alarms

 

PIQUET

  • Discussion of alarms week 04:
  • ALICE TRD Buffer pressure too high due to atmospheric pressure (but also because Chilo changed the pressure in the loop and alarms limit were not changed)
  • CMS RPC: Pump regulation due to increase of the flow —> need to regulate with manual valve in the pump
  • LHCb RICH2: PLC crash. PLC will be send to Schiender for diagnostic. Remember to call BE piquet when crash happen and do not intervene by yourself
    • Patrick will prepare a box with all spare material for PLC. Box will be in electrical workshop
    • For system with old PLCs: Michal to add in the rack an old spare PLC
    • For TFC: Michal to ask to Majec for a spare IO card (special ne for TFC)

 

GENERAL

  • PLC problem of crashes
    • 170 PLCs installed at CERN
    • In total 10 crashes of PLCs at CERN, gas team has 3 crashes —> Michal to crosscheck with Marc about the numbers
    • Schiender said in general that grounding is not good (leak of current)
    • In cryogenic also found some bugs in firmware after the crashes
    • Slide on the modifications asked by Schniedr with pictures and time needed to do modification
  • Beatrice asked to LPh, Patrick, Kacper, Andrea to check how many sensors to install in the cavern: now or in the future. Beatrice will take care of ordering fro FirstSensor the special series for radiation —> by week 7

From previous meeting

  • To fix the StepWS problem asap otherwise difficult to have history in case of problems —> Michal/Gianluca
  • MFC exhaust problem: software or hardware —> Roberto to try with CMS CSC and Michal to investigate
    • CMS DT problem in Exhaust MFC: if you put open it goes in error and you have to disconnect
    • Everybody to check that range of MFC in software is the same of hardware
  • Aging test priority
    • To do: Electrovalve dirty and cleaned of analysis, Vogtlin flowmeter, IR BINOS (they do aging —> LPh to check for modification)

 

 

 

There are minutes attached to this event. Show them.
    • 14:00 14:05
      Piquet interventions 5m

      http://eloggas.cern.ch/GAS_PIQUET/

      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Kacper Jerzy Kapusniak (CERN), Michal Zbigniew Zimny (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR))
    • 14:05 14:15
      M&O and project organization 10m
      Speakers: Beatrice Mandelli (CERN), Roberto Guida (CERN)
    • 14:15 14:25
      Overview of planning LS2 10m
      Speaker: Beatrice Mandelli (CERN)
    • 14:25 14:30
      Activities on-going 5m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN)
    • 14:30 14:50
      Plan of LS2 activities 20m
      Speakers: Beatrice Mandelli (CERN), Frederic Merlet (CERN), Roberto Guida (CERN)
    • 14:50 16:15
      Gas systems for the LHC experiments
      • 15:05
        ALICE 5m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Louis-Philippe De Menezes (CERN)
      • 15:10
        ATLAS 5m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Kacper Jerzy Kapusniak (CERN)
      • 15:15
        CMS 5m
        Speakers: Andrea D'Auria (CERN), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 15:20
        LHCb 5m
        Speakers: Mara Corbetta (Universite Claude Bernard Lyon I (FR)), Patrick Carrie (CERN)
      • 15:25
        Not-LHC experiments 5m
        Speaker: Fabio Bordon (CERN)
      • 15:30
        Software 10m
        Speakers: Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Michal Zbigniew Zimny (CERN)
      • 15:40
        Gas web-page 5m

        http://test-gasmonitoring.web.cern.ch
        https://ep-dt-gasanalysis.web.cern.ch

        Speaker: Beatrice Mandelli (CERN)
      • 15:45
        Gas analysis: standard H2O/O2 5m
        Speaker: Louis-Philippe De Menezes (CERN)
    • 16:15 17:05
      Gas System developments
      • 16:15
        Gas System Test-Bench 10m
        Speaker: Fabio Bordon (CERN)
      • 16:25
        CMS CF4 recovery 10m
        Speakers: Fabio Bordon (CERN), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 16:35
        LHCb RICH2 CF4 recuperation 10m
        Speaker: Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 16:45
        R134a Recuperation 10m
        Speakers: Federico Consolaro (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Ian Crotty (University of Wisconsin Madison (US)), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
      • 16:55
        AoB 10m
        Speakers: Fabio Bordon (CERN), Gianluca Rigoletti (Universite Claude Bernard Lyon I (FR)), Leonardo Nesti (Universita Di Bologna (IT)), Mara Corbetta (Universite Claude Bernard Lyon I (FR))
    • 17:05 17:15
      FSU & Jobs status 10m
      Speaker: Frederic Merlet (CERN)
    • 17:15 17:25
      AoB 10m
      Speakers: Andrea D'Auria (CERN), Beatrice Mandelli (CERN), Frederic Merlet (CERN), Louis-Philippe De Menezes (CERN), Patrick Carrie (CERN), Roberto Guida (CERN)